OpenStudy (anonymous):

By rationalizing the denominators, simplify (12*sqrt(5))/(2*srqt(5)-4)

OpenStudy (anonymous):

\[ \frac{ 12\sqrt 5 }{ 2\sqrt 5-4 }\]

OpenStudy (anonymous):

To do this, you have to multiply the numerator & denominator by the "conjugate" of the denominator.

OpenStudy (anonymous):

In this case, the conjugate is \[2\sqrt{5}+4\]

OpenStudy (anonymous):

When you are looking at a denominator that is a binomial with a square root, that part is critical as any other root involves a much more rigorous process.

OpenStudy (anonymous):

Remember, \[(x+y)*(x-y) = x^{2}-y ^{2}\]

OpenStudy (anonymous):

So you denominator is (\[2^{2} * \sqrt{5}^{2} - 4^{2}\]

OpenStudy (anonymous):

2^2=4 \[\sqrt{5}*\sqrt{5}=5\] -(4^2) = -(16) = -16

OpenStudy (anonymous):

4 * 5 -16 = 4

OpenStudy (anonymous):

Now for the numerator.

OpenStudy (anonymous):

\[(12\sqrt{5})(2\sqrt{5}-4)=\]

OpenStudy (anonymous):

\[(12\sqrt{5}) (2\sqrt{5}-4) = [(12 * 2) (\sqrt{5} * \sqrt{5})] - [(12 * 4) * \sqrt{5)}] = [(24 * 5)] - [(48\sqrt{5})] = 120 - 48\sqrt{5}\]

OpenStudy (anonymous):

\[(12\sqrt{5}) * (2\sqrt{5} -4) =\] \[(12*2)*(\sqrt{5}*\sqrt{5}) - ((12*4)*\sqrt{5})) =\] \[(24) * (5) - (48 * \sqrt{5}) =\] \[120 - 48\sqrt{5}\]

OpenStudy (anonymous):

The final/simplified/rationalized answer is: \[(120 - 48\sqrt{5}) / 4\]